OpenStudy (anonymous):

Barton has two fish tanks in the shape of rectangular prisms. Each of the sides of the second fish tank are 4 times greater than the sides of the first fish tank. If Barton fills both the fish tanks completely with water, the volume of water in the second fish tank will be 16 times the volume of water in the first fish tank 8 times the volume of water in the first fish tank 4 times the volume of water in the first fish tank 64 times the volume of water in the first fish tank

OpenStudy (anonymous):

help please!!!

OpenStudy (anonymous):

volume is given by the equation V=a*b*c. Volume for the first fish tank is V1=x*x*x=x³. Volume for the second tank is V2=4x*4x*4x=4³*x³ comparing the two volumes, we get \[V1/V2 = (x^3)/(64*x³)\] we can cancel out the x³, and then it's easy to see that V2 = 64*V1